OpenStudy (anonymous):

For all x > 0 and y > 0, the radical expression: √x divided by 3√x-√y is equivalent to. 3x + √xy divided by 9x-y. How did they get this answer?

OpenStudy (anonymous):

they just multiplied and divided by 3√x+√y

OpenStudy (mathteacher1729):

When you have the sum or differences of two square roots in your denominator, generally you "multiply by the conjugate" to eliminate them. so if you have say, 1 / ( sqrt(a) + sqrt(b) ) you would multiply THAT by the fraction sqrt(a) - sqrt(b) over sqrt(a) - sqrt(b) (note the minuses!) The reverse is true if you have say, 1 / ( sqrt(a) - sqrt(b) ) you would multiply THAT by the fraction sqrt(a) + sqrt(b) over sqrt(a) + sqrt(b)

OpenStudy (anonymous):

yea conjugate :D

OpenStudy (anonymous):

I'm sorry but I still don't understand.

OpenStudy (anonymous):

Where did the rest of the numbers come from?

OpenStudy (anonymous):

do u know, (a-b)(a+b) = a^2 - b^2

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well, it is. They hv used it. √x / 3√x-√y = (√x )* (3√x+√y) / (3√x-√y )* (3√x+√y) = (3x + √xy) / (9x-y)

OpenStudy (anonymous):

Oooh ok thank you very much....i understand now!

OpenStudy (anonymous):

You're welcome! :)